OpenStudy (anonymous):

major help please, do not understand at all Divide 3x2 – 11x – 4 by x – 4. x + 1 x – 23 3x + 1 3x – 23

OpenStudy (anonymous):

3x2−11x−4=(x−4)(3x+1) Does this help?

OpenStudy (anonymous):

Eh, could you elaborate more?

OpenStudy (anonymous):

What he means is to cancel off the (x-4) term leaving only (3x+1) remaining

OpenStudy (anonymous):

So, I do 3x2-11x-4=(3x+1)

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Good luck with your studies in the Gospel of Mathematics @Zale101 ..I wish you the best of luck! =)

OpenStudy (anonymous):

-8x^2-4=-12(3x+1)=48

OpenStudy (anonymous):

-48